OpenStudy (anonymous):

Factor completely 5c^5 + 60c^4 + 180c^3. 5c^3(c + 6)(c - 6) 5c^3(c^2 + 12c + 36) 5c^3(c + 6)^2 5(c + 6)^2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I'm not sure what the GCF would be on this one

ganeshie8 (ganeshie8):

first look at numbers

ganeshie8 (ganeshie8):

5 , 60, 180 wats the gcf of these ?

OpenStudy (anonymous):

I want to say 5

ganeshie8 (ganeshie8):

yup! pull that out first

OpenStudy (anonymous):

because anything over 5 couldn't be a factor of 5 lol

ganeshie8 (ganeshie8):

5c^5 + 60c^4 + 180c^3 5(c^5 + 12c^4 + 36c^3)

ganeshie8 (ganeshie8):

fine ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

good, next look at c's wats the GCF ?

OpenStudy (anonymous):

c^3?

ganeshie8 (ganeshie8):

perfect ! pull that out also wat do u get ?

OpenStudy (anonymous):

5c^3(c^2 + 12c^1+ 36)

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

:D

ganeshie8 (ganeshie8):

5c^3(c^2 + 12c^1+ 36) is same as 5c^3(c^2 + 12c+ 36)

ganeshie8 (ganeshie8):

we're half done. next, do u knw (a+b)^2 formula ?

OpenStudy (anonymous):

umm not really

ganeshie8 (ganeshie8):

google and tell me :)

OpenStudy (anonymous):

umm expanding formula?

ganeshie8 (ganeshie8):

yup

OpenStudy (anonymous):

couldn't find much on it

ganeshie8 (ganeshie8):

here it is, \(\large (a+b)^2 = a^2 + 2ab + b^2\)

ganeshie8 (ganeshie8):

now look at wat we have :- 5c^3(c^2 + 12c+ 36)

ganeshie8 (ganeshie8):

how can we use that formula here ?

OpenStudy (anonymous):

uhh I'm not sure which one would be a or b if theres 3 numbers

ganeshie8 (ganeshie8):

let me show u how to do this completely :- \(\large 5c^3(c^2 + 12c+ 36)\) \(\large 5c^3(c^2 + 2*6*c+ 6^2)\) \(\large 5c^3(c+6)^2\)

ganeshie8 (ganeshie8):

we're done. see if that makes sense

OpenStudy (anonymous):

Yeah I think I see it

ganeshie8 (ganeshie8):

good :) below 3 identites are very useful, you better remember these for the rest of the course time :- \(\large 1)~~~ (a+b)^2 = a^2+2ab + b^2 \) \(\large 2)~~~ (a-b)^2 = a^2-2ab + b^2\) \(\large 3)~~~ a^2-b^2 = (a+b)(a-b)\)

ganeshie8 (ganeshie8):

oly 3 formulas, take time to memorize these. you should be able to call these at will and apply

OpenStudy (anonymous):

Alright thanks man :)

ganeshie8 (ganeshie8):

np :)